# Argumentative Experience: Reducing Confirmation Bias on Controversial Issues through LLM-Generated Multi-Persona Debates

### v2

@@ 6.2. Balancing Divergent Viewpoints and Convergent Solutions
While participants appreciated the system for broadening their understanding of controversial topics through multiple viewpoints, many expressed a desire for more concrete solutions to address the social issues being debated.
For example, some participants preferred concrete answers over the exchange of opinions. P13 noted personas“tend to stop at potential angles for the topic, but they don’t dive much deeper into more specific details.” P29 said “the chat didn’t address specific concerns and I couldn’t find a clear solution.” This led us to further reflect on when a system should offer diverse perspectives vs. more concrete answers. Providing convergent viewpoints can help spur decision-making but may also lead to automation bias in users too readily accepting system answers, something we sought specifically to avoid in our design. Bearing a more critical and constructive stance, we consider there are different design opportunities to balance divergent viewpoints with convergent solutions effectively.
@@ 6.3. Addressing Potential Bias in LLM-generated Personas
LLM simulation may exhibit many inaccuracies vs. reality due to hallucination, random errors, and consistent biases. Approximation of diversity in the real world may over- or under-representing certain groups and perspectives. Such bias often stems from the limitations of training data that incorporate existing social inequalities, stereotypes, or incomplete views of the population (Christian, 2020). Recent work has examined these misalignment across misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
@@
Of course, users may hold their own biases toward certain social roles which could further reinforce their confirmation biases. The very flexibility our system enables users to customize the LLM personas and correct biases they find also means that they have the power to customize personas that reaffirm their existing beliefs. We cannot objectively evaluate how well participant customization reflects social roles within the demographic groups embodied by those personas. Future work should involve systematic socio-demographic studies to better understand, align, and compare user interpretations of personas with the biases present in LLM-generated personas (Kirk et al., 2024). Additionally, HCI research might consider how to balance giving users control over customization with ensuring representational diversity and fairness in LLM-generated personas.
6.4. Reducing Automation Bias in AI-Assisted Decision-Making
Although we did not explicitly assess the system’s impact on automation bias (Bertrand et al., 2022) (i.e., whether participants would automatically follow AI suggestions or engage in deliberate decision-making), the debate format of arguing opposing positions directly avoids providing simple answers to users that might be accepted without deliberate engagement. The debate format, with high customization, forces users to think critically and objectively assess the different arguments presented. This contrasts with the more passive role users typically take in traditional AI-assisted decision-making scenarios, where they might simply accept AI decisions (Buçinca et al., 2021; Lai et al., 2023).
Many participants intentionally structured a robust debate by selecting or customizing personas with different stances on the question (P14, 22, 24). P22 stated, “I felt like the descriptions that it gave would give a good debate” P24 added, “I’d probably change these personas a little more into what I feel like would really structure a good debate.” Although asking people to grapple with uncertainty may be uncomfortable, this encourages users to engage with multiple viewpoints and think critically to promote more deliberate decision-making. We were inspired in part by prior work using cognitive forcing functions to mitigate user over-reliance on AI decisions (Buçinca et al., 2021; Gajos and Mamykina, 2022). The debate system also offers this opportunity to help provide potential opposing viewpoints, fostering appropriate reliance in AI-assisted group decision-making (Chiang et al., 2024).
@@ 6.5.1. System design and development
As described in Section 6.1, our current multi-persona debate system lacks some interactive features to increase user engagement, such as the ability to nudge debate conversations by incorporating user feedback and new inputs. Additionally, system improvements are needed to enhance debate quality (Section 6.2), such as integrating RAG to reduce potential hallucination problems, refining prompting techniques to minimize repetitive responses from personas, and facilitating a natural closure in the debate (e.g., “I have nothing new to add”).
@@ 40 Female 20 High school or equivalent Applied sciences
Appendix E Additional Findings regarding User Behavior Patterns
Beyond the main behavior clusters outlined in Section 5.1, in this section, we provide additional analysis, describing the topical differences within these groups (Section E.1) and additional qualitative findings regarding how our participants select and revise personas (Section E.2).
@@ E.1. Topical differences among user behavioral clusters
E.2. Persona expertise, individuality, and diversity in selecting and revising personas
From the post-task interviews, we found three additional factors that participants frequently mentioned in selecting and revising personas, including persona expertise, individuality, and the overall diversity of personas when selecting and editing them to construct the debate. These factors were closely tied to how participants conceptualized the controversial topics being discussed.
First, many participants preferred selecting both expert and individual personas. Expert personas were more trusted to provide factual information, evidence, and other logical and reliable arguments to the controversial topic (P11, 14-15, 22, 24, 33, 36, 38-39). P39 explained that for the topic of “TikTok ban,” experts could provide opinions rather than emotional responses “Parents who are concerned about their kid is great, but I don’t expect them to have all of the data necessary to make an informed decision. I thought there would be experts discussing things logically and providing evidence-based information, as opposed to information driven by emotions.” In contrast, individual personas could provide direct and authentic arguments (P11, 22, 24-25, 33, 38). For example, P38 said “Whether TikTok should be banned from school. I really want from people who are in the school specifically because it’s important to see the people who are actually being affected to know their pain points. So that’s why I picked the teacher, the student, and the administrator.
Additionally, as participants were aware that potential bias could underlie different personas, participants aimed to create a balanced representation of opinions by selecting personas that supported different sides of the topic. On one hand, this approach could be an effective strategy for generating counterarguments (P14, 24, 26, 28, 33, 35-36), as P36 explained, “I tried to balance the number of people who supported the topic with those who opposed it, to sort of weigh which perspective truly aligns with my own thinking and against it.” On the other hand, some would deliberately balance opinions that might be disproportionately represented on one side in real-world discussions (P22, 24, 33). P33 pointed out that“If I’ve been hearing a lot of one point of view, I’ll try to keep generating a persona until I find one that opposes it.
Appendix F Additional Statistical Results
F.1. Descriptive user behavior data in debate system
@@
F.2. Generalized linear mixed-effects model analysis
@@
F.3. Distribution of bi-directional belief changes across topics
@@
Table 9. Bi-directional belief changes across different topics: The retrieval-based and multi-persona systems have varying effects on user belief changes depending on the topic. For topics 1 and 2, the retrieval-based system tends to reinforce participants’ previous stances, making them more supportive (highlighted as orange). In contrast, for topics 5 and 8, the multi-persona debate system leads participants to become less supportive of their initial stance (highlighted as lightblue). Both systems only show subtle impacts on other topics, but they both lead to participants becoming more supportive on topic 6 (highlighted as gray).
By mapping participant bidirectional belief changes across different topics (Table 9), we found that the yielded system effects on weighted belief changes also vary on the topic. For example, in topics “Should animals be used for scientific or commercial testing?” and “Should humans colonize space?” (highlighted as orange) participants became more supportive of their previous stances after using the baseline. In contrast, for topics “Should the Federal corporate income tax rate be raised?” and “Do violent video games contribute to youth violence?” (highlighted as blue) participants became less supportive of their previous stance after using the experiment. This suggests that user belief changes differ depending on the topic, with some topics being particularly sensitive.

### v3

@@ 6.2. Balancing Divergent Viewpoints and Convergent Solutions
While participants appreciated the system for broadening their understanding of controversial topics through multiple viewpoints, many expressed a desire for concrete solutions to address the social issues being debated. Thus, in this section, we reflect on participants’ qualitative feedback and explore how to extend the technology-facilitated critical thinking tool beyond viewpoint exploration toward a solution-oriented space.
During our interview, some participants preferred concrete answers over the exchange of opinions. P13 noted personas“tend to stop at potential angles for the topic, but they don’t dive much deeper into more specific details.” P29 said “the chat didn’t address specific concerns and I couldn’t find a clear solution.” This preference for concrete solutions may stem from participants’ increased exposure to opposing viewpoints, which could heighten cognitive dissonance. While no participants explicitly mentioned experiencing cognitive dissonance, such as complaints about divergent viewpoints, their repeated inquiries about clear resolutions might suggest an underlying discomfort with unresolved tensions in the discussion. These observations led us to further reflect on when a system should offer diverse perspectives vs. more concrete answers. Note that providing convergent viewpoints can help spur decision-making but may also lead to automation bias in users too readily accepting system answers, something we sought specifically to avoid in our design. Bearing a more critical and constructive stance, we consider there are different design opportunities to balance divergent viewpoints with convergent solutions effectively.
@@ 6.3. Addressing Potential Bias in LLM-generated Personas
LLM simulation may exhibit many inaccuracies vs. reality due to hallucination, random errors, and consistent biases. Approximation of diversity in the real world may over- or under-representing certain groups and perspectives. Such bias often stems from the limitations of training data that incorporate existing social inequalities, stereotypes, or incomplete views of the population (Christian, 2020). Recent work has examined these misalignments across misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
@@
Of course, users may hold their own biases toward certain social roles, which could further reinforce their confirmation biases. The flexibility of our system enables users to customize the LLM personas and correct biases they find, which also means that they have the power to customize personas that reaffirm their existing beliefs. We cannot objectively evaluate how well participant customization reflects social roles within the demographic groups embodied by those personas. Future work should involve systematic socio-demographic studies to better understand, align, and compare user interpretations of personas with the biases present in LLM-generated personas (Kirk et al., 2024). Additionally, HCI research might consider how to balance giving users control over customization with ensuring representational diversity and fairness in LLM-generated personas.
6.4. Addressing Automation Bias in AI-Assisted Decision-Making
While our study primarily examines the use of LLM multi-personas in reducing confirmation bias, other studies have explored similar applications in the context of AI-assisted decision-making. For example, Chiang et al. (2024) found that LLM could play as a devil advocate to foster user appropriate reliance on AI decisions. This suggests that LLM-driven persona debates may also help mitigate automation bias (Bertrand et al., 2022) (i.e., whether participants would automatically follow AI suggestions or engage in deliberate decision-making). In this section, we discuss relevant insights from participant feedback on this potential benefit.
The debate format of arguing opposing positions directly avoids providing simple answers to users that might be accepted without deliberate engagement. Its high level of customization encourage users to think critically and objectively assess the different arguments presented. This contrasts with the more passive role users typically take in traditional AI-assisted decision-making scenarios, where they might simply accept AI decisions (Buçinca et al., 2021; Lai et al., 2023).
For example, many participants intentionally structured a robust debate by selecting or customizing personas with different stances on the question (P14, 22, 24). P22 stated, “I felt like the descriptions that it gave would give a good debate” P24 added, “I’d probably change these personas a little more into what I feel like would really structure a good debate.” Although asking people to grapple with uncertainty may be uncomfortable, this encourages users to engage with multiple viewpoints and think critically to promote more deliberate decision-making. We were inspired in part by prior work using cognitive forcing functions to mitigate user over-reliance on AI decisions (Buçinca et al., 2021; Gajos and Mamykina, 2022). The debate system also offers this opportunity to help provide potential opposing viewpoints, fostering appropriate reliance in AI-assisted group decision-making (Chiang et al., 2024).
@@ 6.5.1. System design and development
Our current multi-persona debate system lacks some interactive features to increase user engagement, such as the ability to nudge debate conversations by incorporating user feedback and new inputs. Additionally, system improvements are needed to enhance debate quality (Section 6.2), such as integrating RAG to reduce potential hallucination problems, refining prompting techniques to minimize repetitive responses from personas, and facilitating a natural closure in the debate (e.g., “I have nothing new to add”).
While our system was built on top of ArgumentSearch to enable viewpoint diversity and did not incorporate traditional search engines as another baseline for comparison given the burden of evidence of confirmation bias observed (Section 2.2), it is important to acknowledge that this choice may limit the generalizability of our findings to more typical search scenarios. Future studies could include comparative evaluations with traditional or commercial search engines to further assess external validity and robustness.
@@ 40 Female 20 High school or equivalent Applied sciences
Appendix E Additional Findings
Beyond the main behavior clusters and qualitative themes outlined in Section 5.1, we present additional quantitative findings. These include participants’ behavioral patterns across different topics (Section E.1) their descriptive behavior data (Section E.2), detailed generalized linear mixed-effects model analysis (Section E.3), and user distribution of bi-directional belief change across topics (Section E.4).
@@ E.1. Topical differences among user behavioral clusters
[…]
Distribution of user behavior patterns across different topics
@@
E.2. Descriptive user behavior data in debate system
@@
E.3. Generalized linear mixed-effects model analysis
@@
E.4. Distribution of directional belief changes across topics
@@
Table 9. Directional belief changes across different topics: The retrieval-based and multi-persona systems have varying effects on user belief changes depending on the topic. For topics 1 and 2, the retrieval-based system tends to reinforce participants’ supportive stance (highlighted as orange). In contrast, for topics 5 and 8, the multi-persona debate system leads participants to become less supportive (highlighted as lightblue). Both systems only show subtle impacts on other topics, but they both lead to participants becoming more supportive on topic 6 (highlighted as gray).
By mapping participant bidirectional belief changes across different topics (Table 9), we found that the yielded system effects on weighted belief changes also vary on the topic. For example, in topics ‘Animal Testing” and ‘Space Colonization’ (highlighted as orange) participants became more supportive of their previous stances after using the baseline. In contrast, for topics ‘Federal Corporate Income Tax Raise’ and ‘Violent Video Games Effect on Youth Violence’ (highlighted as blue) participants became less supportive of their previous stance after using the experiment. This suggests that user belief changes differ depending on the topic, with some topics being particularly sensitive.
Reflecting on user qualitative insights as summarized in Section 5.1.2 and 5.1.3, we learned that topics like ‘Animal Testing’ and ‘Space Colonization’ may have been influenced by the retrieval system’s ability to present factual, research-driven information, reinforcing expert arguments in favor of these issues. In contrast, the interactive format of the debate system may have heightened participant exposure to counterarguments, leading to decreased support for policies that heavily influence each individual, such as ‘Federal Corporate Income Tax Raise’ and ‘Violent Video Games Effect on Youth Violence.’ For these topics, opposition arguments presented by the debate system tend to be more persuasive. Moreover, the overall increase in support for ‘Universal Basic Income’ across both systems suggests a general receptivity to economic policies that emphasize social welfare.